Part of $18,000 is invested at 8% annual interest, and the rest is invested at 6%. After one year, the accounts paid $1,220 in interest. How much was invested at the lower rate? (Give your answer correct to the nearest cent.)

Answer by nerdybill(7384)

Part of $18,000 is invested at 8% annual interest, and the rest is invested at 6%. After one year, the accounts paid $1,220 in interest. How much was invested at the lower rate? (Give your answer correct to the nearest cent.)
.
Let x = amount invested at lower rate of 6%
then
18000-x = amount invested at 8%
.
Our equation:
.06x + .08(18000-x) = 1220
.06x + 1440 - .08x = 1220
1440 - .02x = 1220
-.02x = -220
x = $11,000.00